# Bayesian Self-Distillation for Image Classification

### v1

@@ A.3 Penultimate Layer Representations
Inspired by muller2019does, we visualize the penultimate layer representations in Figure 13. BSD yields tighter, less overlapping clusters than conventional training, which is somewhat surprising since BSD promotes learning of sample-specific features. It seems that learning similarities between classes can help differentiate among them.
@@ A.4 Limitations
While we experiment with various forms of data augmentation, the interaction with different augmentation schemes as well as regularization techniques warrants further study. Intuitively, augmentations that increase prediction variance may benefit from higher values of the discount factor $\gamma$. Finally, BSD requires selecting the discount factor $\gamma$ and the prior strength $c$, which, despite the observed performance across a large range of settings, could be viewed as a methodological limitation.

### v2

@@ A.3 Penultimate Layer Representations
Inspired by Müller et al. (2019), we visualize the penultimate layer representations in Figure 12. DPS yields tighter, less overlapping clusters than conventional training, which is somewhat surprising since DPS promotes learning of sample-specific features. It seems that learning similarities between classes can help differentiate among them.
@@ A.4 Limitations
While we experiment with various forms of data augmentation, the interaction with different augmentation schemes as well as regularization techniques warrants further study. Intuitively, augmentations that increase prediction variance may benefit from higher values of the discount factor $\gamma$. Finally, DPS requires selecting the discount factor $\gamma$ and the prior strength $c$, which, despite the observed performance across a wide range of settings, could be viewed as a methodological limitation.